OpenStudy (anonymous):

What value of h makes the statement true? (17 + h) + (–9) = 17 + (5 + (–9)) A. –21 B. –5 C. 5 D. 21

OpenStudy (shinalcantara):

(17+h)+(-9) = 17+(5+(-9)) solve first the innermost parenthesis of the right side making it (17+h)+(-9) = 17+(5-9) (17+h)+(-9) = 17+(-4) (17+h)+(-9) = 17-4 (17+h)+(-9) = 13 add both sides of the equation with 9 so that -9 of the left side will be eliminated or you just have to transpose it (17+h)+(-9+9) = 13+9 17+h=22 subtract 17 to both sides of the equation -17+17+h=22-17 h=5

OpenStudy (anonymous):

thank you so much!!!!!!!!!!

OpenStudy (shinalcantara):

yw